# Entropy of self-avoiding branching polymers: mean field theory and Monte Carlo simulations

### v1

@@ Abstract
where $\{{\mathcal{C}}\}_{{\mathbf{x}}_{0}}$ is the set of all possible conformations of directed lattice tree rooted on site ${\mathbf{x}}_{0}$. The total number of bonds in a given conformation ${\mathcal{C}}$ is indicated with $N_{b}({\mathcal{C}})$ and $\kappa_{b}$ is the corresponding bond fugacity; then, $N_{f}({\mathcal{C}})$ is the number of $f$-functional branch-nodes and $\kappa_{f}$ is the corresponding fugacity ($=e^{\beta\mu_{f}}$, where $\mu_{f}$ is the chemical potential for the $f$-functional branch-nodes, see Eq. (8) in main text).
[…]
@@
    $\langle S_{i}({\mathbf{x}})\rangle=1$, for $i=1,...,n$.
@@
    $\langle S_{i}({\mathbf{x}})\,S_{j}({\mathbf{x}})\rangle=0$, for every $i,j$.
@@
    $\langle S_{i}({\mathbf{x}})\,{\bar{S}}_{j_{1}}({\mathbf{x}})\,{\bar{S}}_{j_{2}}({\mathbf{x}})\,\dots\,{\bar{S}}_{j_{k}}({\mathbf{x}})\rangle=\begin{cases}\delta_{i,j_{1}}&\text{if $k=1$;}\\ \delta_{i,j_{1},j_{2}}[\kappa_{3}\,(1-\delta_{{\mathbf{x}},{\mathbf{x}}_{0}})+\delta_{{\mathbf{x}},{\mathbf{x}}_{0}}]&\text{if $k=2$;}\\ \delta_{i,j_{1},j_{2},j_{3}}[\kappa_{4}\,(1-\delta_{{\mathbf{x}},{\mathbf{x}}_{0}})+\kappa_{3}\,\delta_{{\mathbf{x}},{\mathbf{x}}_{0}}]&\text{if $k=3$;}\\ .\\ .\\ \delta_{i,j_{1},j_{2},\dots,j_{q}}\,\kappa_{q}\,\delta_{{\mathbf{x}},{\mathbf{x}}_{0}}&\text{if $k=q$};\\ 0&\text{otherwise.}\end{cases}$, for every $\mathbf{x}$ and ${\mathbf{x}}^{\prime}$ and where the generalized Kronecker’s symbol $\delta_{ij...}$ is $=1$ for identical subscripts and $=0$ otherwise.
@@
where $\Delta(\mathbb{x},\mathbb{x}^{\prime})$ is the adjacency matrix of the lattice. The limit “$n\to 0$” in Eq. (S2) implements the constraint that all polymer conformations with closed loops have weight $=0$. Then, the term “$1$” corresponds to an empty lattice site, the term “$S_{1}({\mathbf{x}}_{0})$” corresponds to a node in the lattice position of the root ${\mathbf{x}}_{0}$, whereas the term “$\bar{\mathbf{S}}({\mathbf{x}})\cdot{\mathbf{S}}({\mathbf{x}}^{\prime})$” represents a directed bond from a node on the site ${\mathbf{x}}$ to a node on the site ${\mathbf{x}}^{\prime}$.
@@
Refer to captionRefer to captionRefer to captionRefer to $\end{array}$
Figure S1: Self-avoiding tree conformations on the $3\times 3$ square lattice with periodic boundary conditions representing some typical terms of the diagrammatic expansion of Eq. (S2). Lattice sites are numbered from 00 to $8$, the spatial position of the tree root is ${\mathbf{x}}_{0}$ (bottom left corner, in yellow).
S1.1.2 Validation of Eq. (S2)
@@
Fig. S1(c) – Here we show how the trace properties rule out any disconnected conformation from the partition function. In fact, the one in the figure corresponds to the term:
@@
It is easy to see that, after factorization, the expression (S9) contains a factor $\sim\langle\bar{S}_{j}({\mathbf{x}}_{5})\rangle$ which, by the trace property (V), is $=0$ and so the weight of the entire conformation.
@@
Again, after factorization, it is straightforward to see that Eq. (S10) contains a factor $\sim\langle S_{j_{1}}({\mathbf{x}}_{7})S_{j_{2}}({\mathbf{x}}_{7})\rangle$ that is $=0$ because of the trace property (IV).
@@
$\kappa_{b}^{4}\,\langle S_{1}({\mathbf{x}}_{0})\,(\bar{\mathbf{S}}({\mathbf{x}}_{4})\cdot{\mathbf{S}}({\mathbf{x}}_{7}))\,(\bar{\mathbf{S}}({\mathbf{x}}_{7})\cdot{\mathbf{S}}({\mathbf{x}}_{8}))\,(\bar{\mathbf{S}}({\mathbf{x}}_{8})\cdot{\mathbf{S}}({\mathbf{x}}_{5}))\,(\bar{\mathbf{S}}({\mathbf{x}}_{5})\cdot{\mathbf{S}}({\mathbf{x}}_{4}))\rangle\,,$ (S11)
@@
$\kappa_{b}^{4}\,\sum_{\begin{subarray}{c}j_{1},j_{2}\\ j_{3},j_{4}\end{subarray}}\,\langle S_{1}({\mathbf{x}}_{0})\rangle\,\langle\bar{S}_{j_{1}}({\mathbf{x}}_{4})S_{j_{4}}({\mathbf{x}}_{4})\rangle\,\langle\bar{S}_{j_{4}}({\mathbf{x}}_{5})S_{j_{3}}({\mathbf{x}}_{5})\rangle\,\langle\bar{S}_{j_{2}}(\mathbb{x}_{7})S_{j_{1}}({\mathbf{x}}_{7})\rangle\,\langle\bar{S}_{j_{3}}({\mathbf{x}}_{8})S_{j_{2}}({\mathbf{x}}_{8})\rangle\,,$ (S12)
@@
$\kappa_{b}^{4}\,\sum_{\begin{subarray}{c}j_{1},j_{2}\\ j_{3},j_{4}\end{subarray}}\delta_{j_{1},j_{2}}\delta_{j_{2},j_{3}}\delta_{j_{3},j_{4}}\delta_{j_{4},j_{1}}=\kappa_{b}^{4}\,n\,.$ (S13)
The weight of the conformation is thus proportional to $n$ yet, owing to the limit $n\to 0$ (see Eq. (S2)), these conformations are also removed from the partition function.
To summarize, we have shown that Eq. (S2) effectively counts only rooted directed SAT conformations and, importantly, with the expected statistical weight given by the fugacities $\kappa_{b}$ and $\{\kappa_{f}\}$. We can now proceed with the derivation of the field-theoretic form of Eq. (S2).
S1.1.3 Field-theoretic form of Eq. (S2)
@@
where the first equality is justified by the fact that in the expansion of the exponential all terms beyond first order have weight $=0$ (see Sec. S1.1.2 for an explanation of this point). In order to obtain the field-theoretic form of Eq. (S14), we apply a generalization of the standard Hubbard-Stratonovich formula Zinn-Justin (2021) to it and get:
@@
[…]
@@
S1.2 Saddle-point approximation
@@
At this point, one can solve (eventually through numerical methods) Eqs. (S22) and (S23) and plug the solutions for $\varphi$ and $\bar{\varphi}$ back into Eq. (S21) to have the mean field estimate of $\Xi$. The saddle-point equations can also be recast in an equivalent form that will be useful in the next Section. By multiplying both sides of Eq. (S23) by $\bar{\varphi}$ it is easy to see that
$1-\frac{\varphi\bar{\varphi}}{q}=\frac{1}{1+\sqrt{\kappa_{b}}\bar{\varphi}+\kappa_{b}\bar{\varphi}\varphi+\bar{\varphi}\sum_{f=3}^{q}\kappa_{b}^{f/2}\kappa_{f}\frac{\varphi^{f-1}}{(f-1)!}}\,.$ (S24)
@@
$1-\rho=\frac{1}{1+\sqrt{\kappa_{b}}\bar{\varphi}+\kappa_{b}\bar{\varphi}\varphi+\bar{\varphi}\sum_{f=3}^{q}\kappa^{f/2}_{b}\kappa_{f}\frac{\varphi^{f-1}}{(f-1)!}}\,,$ (S28)
and thus, because of Eq. (S24), we have that:
@@
$s^{{\mathbf{x}}_{0}}=\frac{1}{V}\log(\Xi)-\rho\log(\kappa_{b})-\rho\sum_{f=3}^{q}\phi_{f}\log(\kappa_{f})=-\rho+\log(1-\rho)-\rho\log(\kappa_{b})-\rho\sum_{f=3}^{q}\phi_{f}\log(\kappa_{f})\,.$ (S36)
@@
S2 A note on [Wagner, Erdemci-Tandogan, Zandi, J. Phys.: Condensed Matter (2015), Ref. Wagner, Erdemci-Tandogan, and Zandi (2015)]
@@
Essentially, the work of Wagner et al. [^5] Notice that similar ideas can be found in works Lubensky and Isaacson (1979); Zilman and Safran (2002). is based on the idea of introducing apt external fields in the Hamiltonian of the magnet such that the “$n\to 0$”-limit becomes equivalent to considering only conformations of self-avoiding branched polymers. Rephrasing in the same language and notation of Sec. S1, we denote here with $J_{1}({\mathbf{x}})$ and $J_{3}({\mathbf{x}})$ the value of the external fields at lattice site of spatial coordinate $\mathbf{x}$. Then, the formalism by Wagner et al. is equivalent to introducing the following trace properties (to be compared to our rules (I)-(VI) presented in Sec. S1) for the spins:
@@
At the same time, all the other traces involving the external fields $J_{1}$ and $J_{3}$ are set $=0$. In terms of these rules, the Wagner-Erdemci-Tandogan-Zandi grand canonical partition function $\Xi_{\rm WETZ}$ of the system is given by
@@
where all fugacities have been set to $=1$ for simplicity. Now notice that, due to the trace rules, the exponential term in Eq. (S37) can be Taylor-expanded and only terms up to the first order need to be retained, since the rest of higher order terms contribute ${\mathcal{O}}(n)$ and, therefore, disappear in the $n\to 0$ limit.
Figure S2: Illustrative conformation of a branched polymer containing a loop on a $3\times 3$ square lattice. Sites of the lattice are numbered from 00 to $8$.
Similarly to Sec. S1.1.2, every term in the sum can be illustrated with a graphical example. Thus a term $\sim J_{1}({\mathbf{x}})$ corresponds to the presence of an end-point on lattice site ${\mathbf{x}}$, while a term $\sim J_{3}({\mathbf{x}})$ indicates the presence of a branch-node (with functionality $=3$). Then, analogously to our model (see Sec. S1.1.1), a term ${\mathbf{S}}({\mathbf{x}})\cdot{\mathbf{S}}({\mathbf{x}}^{\prime})$ represents an undirected bond between lattice positions ${\mathbf{x}}$ and ${\mathbf{x}}^{\prime}$. Consider now the conformation depicted in Fig. S2, whose weight corresponds to the expression:
@@
S3 Enumeration of spanning trees
With $\rho=1$, Eq. (5) in main text gives the entropy of spanning trees of $V$ nodes and given number of branch-nodes $N_{1},N_{2},...,N_{q}$. Thus, by multiplying by $V$ both sides of Eq. (5) in main text and then exponentiating, we get the mean field estimate of the total number of such trees. In order to obtain the total number of spanning trees (i.e. with unrestricted node functionality), this quantity must be summed over all possible values $N_{1},N_{2},\dots,N_{q}$ with constraints Eqs. (1) and (2) in main text. As, however, this procedures is quite unpractical, we consider the following, and much simpler, alternative. We switch back to the ensemble where the total number of nodes is fixed whereas every $N_{f}$ can fluctuate under control of the appropriate external fugacity $\kappa_{f}$. Similarly to Sec. S1.3, we perform a Legendre transform and, by indicating with ${\tilde{s}}={\tilde{s}}_{\rho}(\kappa_{3},\dots,\kappa_{q})\equiv{\tilde{s}}_{\rho}(\{\kappa_{f}\})$ the entropy per site in this new ensemble, we have:
${\tilde{s}}_{\rho}(\{\kappa_{f}\})=s_{\rho}(\{\phi_{f}\})+\rho\sum_{f=3}^{q}\phi_{f}\log\kappa_{f}$ (S39)
where the $\phi_{f}$’s are functions of the $\kappa_{f}$ via Eq. (S35). Now, the entropy per lattice site of unweighted spanning trees in the thermodynamic limit is simply given by ${\tilde{s}}_{\rho=1}(\kappa_{3}=1,\dots,\kappa_{q}=1)$. In Table S1, we compare this result with some exact values for different lattice architectures reported in Refs. Wu (1977); Shrock and Wu (2000). From the comparison, it is apparent how our MF approach becomes more and more accurate with increasing lattice coordination number $q$.
    $q$     Exact Wu (1977); Shrock and Wu (2000)     Bethe approximation De Los Rios, Lise, and Pelizzola (2001) / Deviation     Mean field (this work) / Deviation

### v2

@@ Abstract
where $\{{\mathcal{C}}\}_{{\mathbf{x}}_{0}}$ is the set of all possible conformations of rooted-directed lattice tree rooted on site ${\mathbf{x}}_{0}$. The total number of bonds in a given conformation $\mathcal{C}$} is indicated with $N_{b}({\mathcal{C}})$ and $\kappa_{b}$ is the corresponding bond fugacity; then, $N_{f}({\mathcal{C}})$ is the number of $f$-functional branch-nodes and $\kappa_{f}$ is the corresponding fugacity ($e^{\beta\mu_{f}}$, where $\mu_{f}$ is the chemical potential for the $f$-functional branch-nodes, see Eq. (8) in main text).
Assign now two $n$-component spin vectors, ${\mathbf{S}}({\mathbf{x}})\equiv(S_{1}({\mathbf{x}}),...,S_{n}({\mathbf{x}}))$ and $\bar{{\mathbf{S}}}({\mathbf{x}})\equiv({\bar{S}}_{1}({\mathbf{x}}),...,{\bar{S}}_{n}({\mathbf{x}}))$, to each lattice site of spatial position $\mathbf{x}$}, and define the trace operation (indicated by $\langle\cdot\rangle$) on the spin vectors characterized by the following mathematical properties:
@@
    $\langle S_{i}({\mathbf{x}})\rangle=1$, for i=1,,ni=1,...,n.
@@
    $\langle S_{i}({\mathbf{x}})\,S_{j}({\mathbf{x}})\text{(anything)}\rangle=0$, for every $i,j$.
@@
    $\langle S_{i}({\mathbf{x}})\,{\bar{S}}_{j_{1}}({\mathbf{x}})\,{\bar{S}}_{j_{2}}({\mathbf{x}})\,\dots\,{\bar{S}}_{j_{k}}({\mathbf{x}})\rangle=\begin{cases}\delta_{i,j_{1}}&\text{if $k=1$};\\ \delta_{i,j_{1},j_{2}}\,[\kappa_{3}\,(1-\delta_{{\mathbf{x}},{\mathbf{x}}_{0}})+\delta_{{\mathbf{x}},{\mathbf{x}}_{0}}]&\text{if $k=2$};\\ \delta_{i,j_{1},j_{2},j_{3}}\,[\kappa_{4}\,(1-\delta_{{\mathbf{x}},{\mathbf{x}}_{0}})+\kappa_{3}\,\delta_{{\mathbf{x}},{\mathbf{x}}_{0}}]&\text{if $k=3$};\\ \vdots\\ \delta_{i,j_{1},j_{2},\dots,j_{q-1}}\,[\kappa_{q}\,(1-\delta_{{\mathbf{x}},{\mathbf{x}}_{0}})+\kappa_{q-1}\,\delta_{{\mathbf{x}},{\mathbf{x}}_{0}}]&\text{if $k=q-1$};\\ \delta_{i,j_{1},j_{2},\dots,j_{q}}\,\kappa_{q}\,\delta_{{\mathbf{x}},{\mathbf{x}}_{0}}&\text{if $k=q$};\\ 0&\text{otherwise.}\end{cases}$,
    for every $\mathbf{x}$ and $\mathbf{x}$$^{\prime}$ and where the generalized Kronecker’s symbol $\delta_{ij...}$ is =1=1 for identical subscripts and =0=0 otherwise.
@@
where $\Delta(\mathbb{x},\mathbb{x}^{\prime})$ is the adjacency matrix of the lattice. Then, the term “11” corresponds to an empty lattice site, the term $S_{1}({\mathbf{x}}_{0})$” corresponds to a node in the lattice position of the root ${\mathbf{x}}_{0}$, whereas the term $\bar{\mathbf{S}}({\mathbf{x}})\cdot{\mathbf{S}}({\mathbf{x}}^{\prime})$” represents a directed bond from a node on the site $\mathbf{x}$} to a node on the site $\mathbf{x}$$^{\prime}$.
@@
[…]
Figure S2: Self-avoiding tree conformations on the $3\times 3$ square lattice with periodic boundary conditions representing some typical terms of the diagrammatic expansion of Eq. (S3). Lattice sites are numbered from 0 to $8$, the spatial position of the tree root is ${\mathbf{x}}_{0}$ (bottom left corner, in yellow).
S2.1.2 Validation of Eq. (S3)
@@
Fig. S2(c) – Here we show how the trace properties rule out almost every (see later) disconnected conformation from the partition function. In fact, the one in the figure corresponds to the term:
@@
It is easy to see that, after factorization, the expression (S10) contains a factor $\sim\langle\bar{S}_{j}({\mathbf{x}}_{5})\rangle$ which, by the trace property (V), is =0=0 and so the weight of the entire conformation.
@@
Again, after factorization, it is straightforward to see that Eq. (S11) contains a factor $\sim\langle S_{j_{1}}({\mathbf{x}}_{7})S_{j_{2}}({\mathbf{x}}_{7})\rangle$ that is =0=0 because of the trace property (IV).
@@
$\kappa_{b}^{6}\,\langle S_{1}({\mathbf{x}}_{0})\,(\bar{\mathbf{S}}({\mathbf{x}}_{0})\cdot{\mathbf{S}}({\mathbf{x}}_{1}))\,(\bar{\mathbf{S}}({\mathbf{x}}_{0})\cdot{\mathbf{S}}({\mathbf{x}}_{3}))\,(\bar{\mathbf{S}}({\mathbf{x}}_{4})\cdot{\mathbf{S}}({\mathbf{x}}_{7}))\,(\bar{\mathbf{S}}({\mathbf{x}}_{7})\cdot{\mathbf{S}}({\mathbf{x}}_{8}))\,(\bar{\mathbf{S}}({\mathbf{x}}_{8})\cdot{\mathbf{S}}({\mathbf{x}}_{5}))\,(\bar{\mathbf{S}}({\mathbf{x}}_{5})\cdot{\mathbf{S}}({\mathbf{x}}_{4}))\rangle\,$, (S12)
@@
$\kappa_{b}^{6}\,\sum_{\begin{subarray}{c}j_{1},j_{2}\\ j_{3},j_{4}\\ j_{5},j_{6}\end{subarray}}\,\langle S_{1}({\mathbf{x}}_{0})\bar{S}_{j_{1}}({\mathbf{x}}_{0})\bar{S}_{j_{2}}({\mathbf{x}}_{0})\rangle\,\langle S_{j_{1}}({\mathbf{x}}_{1})\rangle\,\langle S_{j_{2}}({\mathbf{x}}_{3})\rangle\,\langle\bar{S}_{j_{3}}({\mathbf{x}}_{4})S_{j_{6}}({\mathbf{x}}_{4})\rangle\,\langle\bar{S}_{j_{6}}({\mathbf{x}}_{5})S_{j_{5}}({\mathbf{x}}_{5})\rangle\,\langle\bar{S}_{j_{4}}(\mathbb{x}_{7})S_{j_{3}}({\mathbf{x}}_{7})\rangle\,\langle\bar{S}_{j_{5}}({\mathbf{x}}_{8})S_{j_{4}}({\mathbf{x}}_{8})\rangle\,$, (S13)
@@
$\kappa_{b}^{6}\,\sum_{\begin{subarray}{c}j_{1},j_{2}\end{subarray}}\,\delta_{1,j_{1},j_{2}}\,\sum_{\begin{subarray}{c}j_{3},j_{4}\\ j_{5},j_{6}\end{subarray}}\,\delta_{j_{3},j_{4}}\delta_{j_{4},j_{5}}\delta_{j_{5},j_{6}}\delta_{j_{6},j_{3}}=\kappa_{b}^{6}\,n\,.$ (S14)
The weight of the conformation is thus proportional to $n$ yet, owing to the limit $n\to 0$ (see Eq. (S3)), these conformations are also removed from the partition function. In fact, it is not difficult to realize that the $n\to 0$ limit helps precisely to get rid of the unwanted case of isolated loops.
To summarize, we have shown that Eq. (S3) effectively counts only rooted-directed SAT conformations and, importantly, with the expected statistical weight given by the fugacities $\kappa_{b}$ and $\{\kappa_{f}\}$. We can now proceed with the derivation of the field-theoretic form of Eq. (S3).
S2.1.3 Field-theoretic form of Eq. (S3)
@@
where the first equality is justified by the fact that in the expansion of the exponential all terms beyond first order have weight =0=0 (see Sec. S2.1.2 for an explanation of this point). In order to obtain the field-theoretic form of Eq. (S15), we apply a generalization of the standard Hubbard-Stratonovich formula Zinn-Justin (2021) to it and get:
@@
where we have already taken the limit for $(\det\Delta^{-1})^{n}\to 1$ and $(2\pi i)^{n}\to 1$. At this point, the term inside the $\langle\rangle$ can be factorized, making the calculation of the trace much easier. By expanding again the exponential and using the trace properties (I)-(VI), one finally gets the expression:
@@
S2.1.4 Direct validation of Eq. (S18) by Wick’s theorem
Eq. (S18) can be readily validated by means of (a classical version of) Wick’s theorem Zinn-Justin (2021). To show it, we take the logarithms and write Eq. (S18) as
$\Xi$ $=$ $\lim_{n\to 0}\int\prod_{\mathbf{x}}d\bm{\varphi}(\mathbb{x})\,d\bar{\bm{\varphi}}({\mathbf{x}})\exp\bigg{[}-\sum_{{\mathbf{x}},{\mathbf{x}}^{\prime}}\,\Delta^{-1}({\mathbf{x}},{\mathbf{x}}^{\prime})\,\bar{\bm{\varphi}}({\mathbf{x}})\cdot\bm{\varphi}({\mathbf{x}}^{\prime})\bigg{]}$ (S19)
$\times\bigg{(}1+\sqrt{\kappa_{b}}\,\varphi_{1}({\mathbf{x}}_{0})+\frac{\kappa_{b}}{2}\,\varphi_{1}({\mathbf{x}}_{0})^{2}+\sum_{f=3}^{q}\kappa^{f/2}_{b}\,\kappa_{f}\,\frac{\varphi_{1}^{f}({\mathbf{x}}_{0})}{f!}\bigg{)}$
$\times\prod_{{\mathbf{x}}\neq{\mathbf{x}}_{0}}\bigg{(}1+\sqrt{\kappa_{b}}\,\sum_{\alpha=1}^{n}\bar{\varphi}_{\alpha}({\mathbf{x}})+\kappa_{b}\sum_{\alpha=1}^{n}\bar{\varphi}_{\alpha}({\mathbf{x}})\,\varphi_{\alpha}({\mathbf{x}})+\sum_{\alpha=1}^{n}\bar{\varphi}_{\alpha}({\mathbf{x}})\sum_{f=3}^{q}\kappa^{f/2}_{b}\,\kappa_{f}\,\frac{\varphi_{\alpha}^{f-1}({\mathbf{x}})}{(f-1)!}\bigg{)}\,.$
Then, computing the integrals of the various terms which appear in Eq. (S19) is equivalent to determining the expectation values of polynomials for a multivariate Gaussian distribution with covariance matrix $\Delta^{-1}$. In particular, we are making use of the following form of the theorem Zinn-Justin (2021):
$\langle\varphi_{\alpha_{1}}(\mathbb{x}_{i_{1}})\bar{\varphi}_{\beta_{1}}(\mathbb{x}_{j_{1}})\dots\varphi_{\alpha_{k}}(\mathbb{x}_{i_{k}})\bar{\varphi}_{\beta_{k}}(\mathbb{x}_{j_{k}})\rangle_{\text{Gauss}}=\sum_{\begin{subarray}{c}\text{all permutations $P$}\\ \text{of $\{1,\dots,k\}$}\end{subarray}}\delta_{\alpha_{1}\beta_{P_{1}}}\Delta(\mathbb{x}_{i_{1}},\mathbb{x}_{j_{P_{1}}})\dots\delta_{\alpha_{k}\beta_{P_{k}}}\Delta(\mathbb{x}_{i_{k}},\mathbb{x}_{j_{P_{k}}})\,$, (S20)
with the notation,
$\langle\cdots\rangle_{\text{Gauss}}\equiv(2i\pi)^{Vn}(\det\Delta)^{n}\int\prod_{\mathbf{x}}d\bm{\varphi}(\mathbb{x})\,d\bar{\bm{\varphi}}({\mathbf{x}})(\cdots)\exp\bigg{[}-\sum_{{\mathbf{x}},{\mathbf{x}}^{\prime}}\,\Delta^{-1}({\mathbf{x}},{\mathbf{x}}^{\prime})\,\bar{\bm{\varphi}}({\mathbf{x}})\cdot\bm{\varphi}({\mathbf{x}}^{\prime})\bigg{]}\,$, (S21)
and where all the moments with a non-equal number of $\varphi$ and $\bar{\varphi}$ terms give 0. For instance, with reference to Fig. S2, let us consider
$\bigg{\langle}\sqrt{\kappa_{b}}\,\varphi_{1}({\mathbf{x}}_{0})\bigg{(}\sum_{\alpha_{1}}\bar{\varphi}_{\alpha_{1}}({\mathbf{x}}_{3})\kappa^{3/2}_{b}\kappa_{3}\frac{\varphi^{2}_{\alpha_{1}}({\mathbf{x}}_{3})}{2}\bigg{)}\bigg{(}\sqrt{\kappa_{b}}\sum_{\alpha_{2}}\bar{\varphi}_{\alpha_{2}}({\mathbf{x}}_{4})\bigg{)}\bigg{(}\sqrt{\kappa_{b}}\sum_{\alpha_{3}}\bar{\varphi}_{\alpha_{3}}({\mathbf{x}}_{6})\bigg{)}\bigg{\rangle}_{\text{Gauss}}=\\ \frac{\kappa^{3}_{b}\kappa_{3}}{2}\sum_{\alpha_{1},\alpha_{2},\alpha_{3}}\langle\varphi_{1}({\mathbf{x}}_{0})\bar{\varphi}_{\alpha_{1}}({\mathbf{x}}_{3})\varphi_{\alpha_{1}}({\mathbf{x}}_{3})\bar{\varphi}_{\alpha_{2}}({\mathbf{x}}_{4})\varphi_{\alpha_{1}}({\mathbf{x}}_{3})\bar{\varphi}_{\alpha_{3}}({\mathbf{x}}_{6})\rangle_{\text{Gauss}}\,.$ (S22)
By using Eq. (S20), it is easy to see that in the limit $n\to 0$ Eq. (S22) simplifies to:
$\kappa^{3}_{b}\,\kappa_{3}\,\Delta({\mathbf{x}}_{0},{\mathbf{x}}_{3})\Delta({\mathbf{x}}_{3},{\mathbf{x}}_{4})\Delta({\mathbf{x}}_{3},{\mathbf{x}}_{6})=\kappa^{3}_{b}\,\kappa_{3}\,$, (S23)
which corresponds (see Eq. (S7)) to the configuration depicted in panel (a) of Fig. S2. By following a similar procedure for all the different terms in Eq. (S19), it can be checked that Eq. (S18) indeed counts all (and only) the SAT configurations with the correct weight based on the total number of bonds and of $f$-functional nodes. In particular, a term $\varphi_{\alpha}({\mathbf{x}})$ indicates a bond exiting from site $\mathbf{x}$}, whereas $\bar{\varphi}_{\alpha}({\mathbf{x}})$ indicates a bond entering in site $\mathbf{x}$}.
S2.2 Saddle-point approximation
@@
At this point, one can solve (eventually through numerical methods) Eqs. (S28) and (S29) and plug the solutions for $\varphi$ and $\bar{\varphi}$ back into Eq. (S27) to have the mean-field estimate of $\Xi$. The saddle-point equations can also be recast in an equivalent form that will be useful in the next Section. By multiplying both sides of Eq. (S29) by $\bar{\varphi}$ it is easy to see that
$\left(1-\frac{\varphi\bar{\varphi}}{q}\right)^{-1}=1+\sqrt{\kappa_{b}}\bar{\varphi}+\kappa_{b}\bar{\varphi}\varphi+\bar{\varphi}\sum_{f=3}^{q}\kappa_{b}^{f/2}\kappa_{f}\frac{\varphi^{f-1}}{(f-1)!}\,.$ (S30)
@@
$(1-\rho)^{-1}=1+\sqrt{\kappa_{b}}\bar{\varphi}+\kappa_{b}\bar{\varphi}\varphi+\bar{\varphi}\sum_{f=3}^{q}\kappa^{f/2}_{b}\kappa_{f}\frac{\varphi^{f-1}}{(f-1)!}\,$, (S34)
and, because the r.h.s. terms of Eq. (S30) and Eq. (S34) are identical, we conclude that:
@@
$\frac{s_{\rho}^{{\mathbf{x}}_{0}}(\{\phi_{f}\})}{k_{B}}=\frac{1}{V}\log(\Xi)-\rho\log(\kappa_{b})-\rho\sum_{f=3}^{q}\phi_{f}\log(\kappa_{f})=-\rho-\log(1-\rho)-\rho\log(\kappa_{b})-\rho\sum_{f=3}^{q}\phi_{f}\log(\kappa_{f})\,.$ (S42)
@@
S3 A note on [Wagner, Erdemci-Tandogan, Zandi, J. Phys.: Condensed Matter (2015), Ref. Wagner, Erdemci-Tandogan, and Zandi (2015)]
@@
Essentially, the work of Wagner et al. [^4] Notice that similar ideas can be found in works Lubensky and Isaacson (1979); Zilman and Safran (2002). is based on the idea of introducing apt external fields in the Hamiltonian of the magnet such that the $n\to 0$”-limit becomes equivalent to considering only conformations of self-avoiding branched polymers. Rephrasing in the same language and notation of Sec. S2, we denote here with $J_{1}({\mathbf{x}})$ and $J_{3}({\mathbf{x}})$ the value of the external fields at lattice site of spatial coordinate $\mathbf{x}$. Then, the formalism by Wagner et al. is equivalent to introducing the following trace properties (to be compared to our rules (I)-(VI) presented in Sec. S2) for the spins:
@@
At the same time, all the other traces involving the external fields $J_{1}$ and $J_{3}$ are set =0=0. In terms of these rules, the Wagner-Erdemci-Tandogan-Zandi grand canonical partition function $\Xi_{\rm WETZ}$ of the system is given by
@@
where all fugacities have been set to =1=1 for simplicity. Now notice that, due to the trace rules, the exponential term in Eq. (S43) can be Taylor-expanded and only terms up to the first order need to be retained, since the rest of higher order terms contribute ${\mathcal{O}}(n)$ and, therefore, disappear in the $n\to 0$ limit.
Figure S3: Illustrative conformation of a branched polymer containing a loop on a $3\times 3$ square lattice. Sites of the lattice are numbered from 0 to $8$.
Similarly to Sec. S2.1.2, every term in the sum can be illustrated with a graphical example. Thus a term $\sim J_{1}({\mathbf{x}})$ corresponds to the presence of an end-point on lattice site $\mathbf{x}$}, while a term $\sim J_{3}({\mathbf{x}})$ indicates the presence of a branch-node (with functionality =3=3). Then, analogously to our model (see Sec. S2.1.1), a term ${\mathbf{S}}({\mathbf{x}})\cdot{\mathbf{S}}({\mathbf{x}}^{\prime})$ represents an undirected bond between lattice positions $\mathbf{x}$} and $\mathbf{x}$$^{\prime}$. Consider now the conformation depicted in Fig. S3, whose weight corresponds to the expression:
@@
S4 Enumeration of spanning trees
With $\rho=1$, Eq. (5) in main text gives the entropy of spanning trees of $V$ nodes and given number of branch-nodes $N_{1},N_{2},...,N_{q}$. Thus, by multiplying by $V$ both sides of Eq. (5) in main text and then exponentiating, we get the mean-field estimate of the total number of such trees. In order to obtain the total number of spanning trees (i.e. with unrestricted node functionality), this quantity must be summed over all possible values $N_{1},N_{2},\dots,N_{q}$ with constraints Eqs. (1) and (2) in main text.
As, however, this procedures is quite unpractical, we consider the following, and much simpler, alternative. We switch back to the ensemble where the total number of nodes is fixed whereas every $N_{f}$ can fluctuate under control of the appropriate external fugacity $\kappa_{f}$. Similarly to Sec. S2.3, we perform a Legendre transform and, by indicating with ${\tilde{s}}={\tilde{s}}_{\rho}(\kappa_{3},\dots,\kappa_{q})\equiv{\tilde{s}}_{\rho}(\{\kappa_{f}\})$ the entropy per site in this new ensemble, we have the simple expression (see also Eqs. (S39) and (S42)):
$\frac{{\tilde{s}}_{\rho}(\{\kappa_{f}\})}{k_{B}}=\frac{s_{\rho}(\{\phi_{f}\})}{k_{B}}+\rho\sum_{f=3}^{q}\phi_{f}\log(\kappa_{f})=\rho\log\left(\frac{q}{e}\right)-(1-\rho)\log(1-\rho)-\rho\log(\phi_{2})\,$, (S45)
where the $\phi_{f}$’s are functions of the $\kappa_{f}$ via Eq. (S41). Now, the entropy per lattice site of unweighted spanning trees in the thermodynamic limit is simply given by
$\frac{{\tilde{s}}_{\rho=1}(\kappa_{3}=1,\dots,\kappa_{q}=1)}{k_{B}}=\log\left(\frac{q}{e}\right)-\log(\phi_{2})\,.$ (S46)
In Table S1, we compare this result with some exact values for different lattice architectures reported in Refs. Wu (1977); Shrock and Wu (2000). From the comparison, it is apparent how our MF approach becomes more and more accurate with increasing lattice coordination number $q$.
    $q$     Exact Wu (1977); Shrock and Wu (2000)     Bethe approximation De Los Rios, Lise, and Pelizzola (2001) / Deviation     Mean-field (this work) / Deviation